# Understanding The Concerns and Choices of The Public When Using Large Language Models for Healthcare

### v1

@@ Abstract.
Large language models (LLMs) have shown their potential in biomedical fields. However, how the public uses them for healthcare purposes such as medical Q&A, self-diagnosis, and daily healthcare information seeking is under-investigated. In this paper, we adopt a mixed-methods approach, including surveys (N=167) and interviews (N=17) to investigate how and why the public uses LLMs for healthcare. LLMs as a healthcare tool have gained popularity, and are often used in combination with other information channels such as search engines and online health communities to optimize information quality. LLMs provide more accurate information and a more convenient interaction/service model compared to traditional channels. LLMs also do a better job of reducing misinformation, especially in daily healthcare questions. Doctors using LLMs for diagnosis is less acceptable than for auxiliary work such as writing medical records. Based on the findings, we reflect on the ethical and effective use of LLMs for healthcare and propose future research directions.
@@ 1. Introduction
Large language models (LLMs) have been a topic of great interest in recent years, especially in the field of natural language processing (NLP). LLMs are a type of AI model designed to generate human-like text by analyzing vast amounts of data. LLMs are based on deep learning techniques and typically involve neural networks with many layers and a large number of parameters, allowing them to capture complex patterns in the data they are trained on (Zhao et al., 2023). The primary goal of LLMs is to understand the structure, syntax, semantics, and context of natural language, so they can generate coherent and contextually appropriate responses. ChatGPT is the first and currently most popular LLM-powered chatbot: as of December 2023, ChatGPT had more than 180 million users (Cha, [n. d.]).
@@
Even before LLMs existed, Internet-based medical services such as healthcare chatbots and ask-the-doctor platforms had gained popularity (Ding et al., 2020). Internet-based self-diagnosis and healthcare information-seeking provide people with convenient and fast medical services without waiting for appointments or queuing, and help people better understand their health status and improve their health awareness. However, Internet-based medical services have such disadvantages as inaccurate information  (Johnson et al., 2022), potentially leading to misdiagnosis or delayed treatment.
Internet-based healthcare information-seeking channels include search engines, online communities, ask-the-doctor platforms, and so on. Now LLMs also provide similar services and potentially bring new risks and opportunities. LLMs’ impact as a healthcare information provider on public health, doctor-patient relationships, and other aspects of people’s personal health has not been well investigated yet. To bridge this research gap, we aspire to answer the following research questions through surveys and interviews:
@@
Through the surveys, we found how the public accesses health care information. and gained a preliminary understanding of the issues that arise when the public obtains health information.
To gain a deeper understanding of the public’s choices and concerns about using LLMs to obtain health information, we conducted semi-structured interviews with 17 participants who had used LLMs in this regard. We found that: (1) people have learned to use LLMs to obtain health information effectively, e.g., through cross-validation with other information channels; (2) compared with search engines, LLMs provide healthcare information more accurately and conveniently; (3) people with medical knowledge are more likely to discover the “illusion” of using LLMs to obtain health information, while for the general public, LLMs provide sufficient information to help them make self-diagnosis; (4) it is controversial whether doctors should use LLMs to diagnose patients and less so when it comes to using LLMs to perform auxiliary work such as writing medical records.
@@
Figure 1. Image of health retrieval using LLMs drawn using Copilot (Microsoft’s large language model). This image is also used for our survey distribution.
@@ 2.1. LLMs in The Healthcare Field
LLMs are advanced AI systems capable of understanding and generating human-like text. They are called “large” because they are trained on vast amounts of data and have a significant number of parameters that help them understand the nuances of language (Zhao et al., 2023). From its development history, LLMs have evolved from transformers (Vaswani et al., 2017) to the current user-facing models such as ChatGPT. Today’s LLMs like ChatGPT (Ray, 2023) and LLaMA (Touvron et al., 2023) are essentially decoder-only transformers (Yang et al., 2023). ChatGPT is the most popular LLM – as of December 2023, ChatGPT had more than 180 million users (Cha, [n. d.]).
The powerful performance of LLMs has attracted the attention of researchers in the healthcare field. Most research focuses on fine-tuning LLMs to better suit medical tasks (Singhal et al., 2023; Liu et al., 2023; Delsoz et al., 2023; Yang et al., 2022). Some studies discuss the application of LLMs in the field of medical education (Sallam, 2023; Safranek et al., 2023). Others have explored the challenges of practical clinical application of LLMs (Harrer, 2023; Singhal et al., 2023; Sezgin, 2023). However, in the field of public healthcare, there are currently few relevant studies. In 2023, a study examined the benefits and challenges of Carecall, an LLM-based conversational robot developed in Korea, for public health applications (Jo et al., 2023). Carecall is aimed at socially isolated individuals, while popular LLMs such as ChatGPT can potentially help the mass public access healthcare information. In this paper, we aim to explore how the public utilizes general-purpose LLMs for healthcare purposes.
Some studies provide suggestions from the perspective of experts on how the public can use LLMs to obtain healthcare information (Biswas, 2023). However, they do not provide an in-depth understanding of the general public’s perspectives, so the proposals given are often from a technical perspective and lack a sociological perspective. A recent study explored the public’s intention to use ChatGPT for self-diagnosis (Shahsavar et al., 2023), but the survey was relatively simple – it only investigated whether the public was willing to use ChatGPT for self-diagnosis, and ignored many other related issues. We instead aspired to provide a nuanced picture of users’ motivations, experiences, and challenges when using LLMs for healthcare.
@@ 2.2. How the Public Access Healthcare Information
The Internet has greatly changed the form of information transmission, and information can be transmitted at an unprecedented speed. Before 2010, the public mainly obtained healthcare information from the Internet through search engines (Berland et al., 2001; Spink et al., 2004; Eysenbach and Kohler, 2003; Gray et al., 2005). After 2010, the methods of obtaining health information gradually diversified. Although search engines were still the most popular channel (De Choudhury et al., 2014), the public also began to try different methods to obtain health information. Some people used social media such as Twitter (De Choudhury et al., 2014; Scanfeld et al., 2010) or WeChat (Zhang et al., 2017; Hong et al., 2021) to obtain or share health information. In China, due to the shortage of medical resources, some online consultation platforms have emerged – patients communicate with doctors online through medical platforms  (Ding et al., 2020; Hao, 2015; Hao et al., 2017). With the rise of artificial intelligence technology, some AI-driven medical/healthcare chatbots have emerged in recent years  (Pereira and Díaz, 2019), such as Ada  (You and Gui, 2020). Patients or the general public can enter questions into the chatbots and get answers. Overall, the ways for the public to obtain health information are becoming increasingly diverse, which is in line with the trend of “information explosion” in the information society. LLMs demonstrate more powerful natural language interaction capabilities and have penetrated the medical field (Li et al., 2023). How the public utilizes them for personal healthcare purposes remains under-investigated.
2.3. Concerns When Accessing Healthcare Information
@@
In the information explosion process, one of the prevalent concerns is that misinformation also explodes (Zhou et al., 2019; Shu et al., 2017). In the field of public health, misinformation is particularly rampant (Suarez-Lledo and Alvarez-Galvez, 2021). Misinformation related to serious diseases (e.g., cancer-related misinformation (Johnson et al., 2022; Chen et al., 2018; Warner et al., 2022)) has often been circulated on social media, accompanied by advertisements for high-priced health products that had no actual therapeutic effect. Since the beginning of the COVID-19 pandemic, conspiracy theories have emerged – for example, rumor went that the COVID-19 virus was spread by 5G technology (Ahmed et al., 2020). A wide range of misinformation about the COVID-19 pandemic existed on social media (Enders et al., 2020; Kouzy et al., 2020; Evanega et al., 2020; Cuan-Baltazar et al., 2020).
@@
Existing studies have pointed out that ChatGPT could provide accurate health information if appropriately prompted (Mehnen et al., 2023). However, considering that LLMs are known to generate “hallucinations” (Alkaissi and McFarlane, 2023; Rawte et al., 2023), understanding how the public experiences “hallucinations” when using LLMs to access healthcare information is important. One study showed that in a Turing test, patients had difficulty distinguishing whether the advice was given by a real doctor or generated by ChatGPT (Nov et al., 2023). Many doctors and scholars have warned that “AI cannot replace doctors (Mehnen et al., 2023; Li, 2023).”
@@
The doctor-patient relationship has always been a controversial topic, which can be complicated by LLMs. As early as 2006, there was research on doctors using search engines to diagnose conditions (Tang and Ng, 2006). In 2011, a survey showed that nearly half of doctors used Google and Yahoo to diagnose patients (Sea, [n. d.]a). In November 2023, a video on the Chinese Internet sparked heated discussions. A doctor diagnosed a patient while swiping the screen of his mobile phone to search. Many Internet users believed that the doctor’s behavior was inappropriate (Sea, [n. d.]b). There has also been a lot of discussion on social media platforms about whether doctors should use search engines to diagnose patients (Goo, [n. d.]). A relevant question is “whether doctors get mad when some patients come with their own Google research and diagnosis” (pat, [n. d.]). These are typical issues and dilemmas in the doctor-patient relationship. Nowadays, the rise of LLMs and their domain knowledge in healthcare and medicine may have an extra impact on the doctor-patient relationship in addition to search engines and online communities. Will patients mind if doctors use LLMs to diagnose them? This question is worth further consideration and research.
@@ 3. Methodology
In this IRB-approved study, we used a combination of interviews and surveys to approach our research questions. Anyone over 18 who had used LLMs for healthcare Q&A or information search was eligible for our study. The two methods are complementary to each other: surveys were administered to understand users’ overall perceptions and ensure the generalizability of the findings; interviews were used to learn about users’ nuanced motivations, experiences, and concerns when using LLMs for health information search.
@@ 3.1. Surveys
Between December 2023 and January 2024, we distributed the online survey on social media platforms such as X, Reedit, and Discord. In the recruitment message, we invited users who satisfied the inclusion criteria to complete our survey. The participation was voluntary, and the participants were offered a chance to enter a prize draw worth 50 dollars (for 7 participants). Since our advertising was primarily distributed through online communities, people who participated in the survey may be limited to those who were familiar with social media.
@@
Interested participants were directed to the survey after clicking on our Qualtrics link. They were first shown information related to the survey; after they gave consent, they proceeded to the survey questions. The survey had a total of 24 questions and the estimated completion time was 10 minutes – in the end, most participants completed the survey within average 5 minutes. When designing survey questions, we referred to previous research on AI healthcare chatbots (Nadarzynski et al., 2019) and took interview results and actual conditions into account. Specifically, the survey was divided into three parts: (1) Demographic information such as gender, race, education status, and so on; (2) How participants searched for healthcare information and what types of healthcare information they searched for – for example, “Have you ever used the search engine (like Google, Bing, Baidu) to search for health information?”, “Have you ever searched for information about medication (treatment, drugs, hospitalization)?” (3) Participants were asked to evaluate controversial issues in the use of LLMs using a 5-point Likert scale – “I think it’s reasonable for doctors to use LLMs to diagnose and treat conditions.”
@@
During the quantitative analysis, we first removed surveys that were duplicates or looked like robot responses. When a large number of consecutive surveys appeared within a short period, we determined them as filled in by robots. We ended up receiving 164 valid surveys. Similar to previous studies, we transformed most of the questions into binary variables and excluded neutral values (Nadarzynski et al., 2019). For race and age, we conducted population statistics by interval. The questions involving value judgments and personal opinions were complex and did not fully conform to the binary distribution. However, they could be used to explore the acceptability of LLMs in public healthcare applications and provide important guidance for future directions.

### v2

@@ Abstract.
Large language models (LLMs) have shown their potential in biomedical fields. However, how the public uses them for healthcare purposes such as medical Q&A, self-diagnosis, and daily healthcare information seeking is under-investigated. This paper adopts a mixed-methods approach, including surveys (N=214) and interviews (N=17) to investigate how and why the public uses LLMs for healthcare. We found that participants generally believed LLMs as a healthcare tool have gained popularity, and are often used in combination with other information channels such as search engines and online health communities to optimize information quality. Based on the findings, we reflect on the ethical and effective use of LLMs for healthcare and propose future research directions.
@@ 1. Introduction
Large language models (LLMs) have been a topic of great interest in recent years, especially in the field of natural language processing (NLP). LLMs are a type of AI model designed to generate human-like text by analyzing vast amounts of data. LLMs are based on deep learning techniques and typically involve neural networks with many layers and a large number of parameters, allowing them to capture complex patterns in the data they are trained on (Zhao et al., 2023). The primary goal of LLMs is to understand the structure, syntax, semantics, and context of natural language, so they can generate coherent and contextually appropriate responses. ChatGPT is the first and currently most popular LLM-powered chatbot: as of December 2023, ChatGPT had more than 180 million users (Duarte, 2024).
@@
Even before LLMs existed, Internet-based healthcare information-seeking channels had gained popularity. Internet-based healthcare information-seeking channels include search engines (De Choudhury et al., 2014; Rice, 2006), online communities (Chung et al., 2017; Loh et al., 2023), ask-the-doctor platforms (Ding et al., 2020; Hao, 2015), healthcare chatbots (You and Gui, 2020; Nadarzynski et al., 2019; Pereira and Díaz, 2019) and so on. Internet-based self-diagnosis and healthcare information-seeking provide people with convenient and fast medical services without waiting for appointments or queuing, and help people better understand their health status and improve their health awareness. However, Internet-based medical services have such disadvantages as inaccurate information  (De Choudhury et al., 2014; Johnson et al., 2022), potentially leading to misdiagnosis or delayed treatment.
Now LLMs also provide similar services and potentially bring new risks and opportunities. LLMs’ impact as a healthcare information provider on public health, doctor-patient relationships, and other aspects of people’s personal health has not been well investigated yet. This is because current research on the ethical issues of the use of LLMs in healthcare mainly focuses on clinical applications and support of health professionals and researchers, and there are few studies on public health (Haltaufderheide and Ranisch, 2024). To bridge this research gap, we aspire to answer the following research questions through surveys and interviews:
@@
Through the surveys, we took a broader perspective to find how the public accesses healthcare information and gained a preliminary understanding of the issues that arise when the public obtains health information.
To gain a deeper understanding of the public’s choices and concerns about using LLMs to obtain health information, we conducted semi-structured interviews with 17 participants who had used LLMs in this regard. We found that: (1) people have learned to use LLMs to obtain health information effectively, and through cross-validation with other information channels; (2) compared with search engines, LLMs provide healthcare information maybe more accurately and conveniently; (3) people with medical knowledge are more likely to discover the “hallucinations” of using LLMs to obtain health information, while for the general public, LLMs provide sufficient information to help them make self-diagnosis; (4) it is controversial whether doctors should use LLMs to diagnose patients and less so when it comes to using LLMs to perform auxiliary work such as writing medical records. (5) In the digital niche, LLMs have both competitive and complementary relationships with search engines and have complementary relationships with online communities and online consultations.
@@ 2.1. LLMs in The Healthcare Field
LLMs are advanced AI systems capable of understanding and generating human-like text. They are called “large” because they are trained on vast amounts of data and have a significant number of parameters that help them understand the nuances of language (Zhao et al., 2023). From its development history, LLMs have evolved from transformers (Vaswani et al., 2017) to the current user-facing models such as ChatGPT. Today’s LLMs like ChatGPT (Ray, 2023) and LLaMA (Touvron et al., 2023) are essentially decoder-only transformers (Yang et al., 2023). ChatGPT is the most popular LLM – as of December 2023, ChatGPT had more than 180 million users (Duarte, 2024).
The powerful performance of LLMs has attracted the attention of researchers in the healthcare field. Most research focuses on fine-tuning LLMs to better suit medical tasks (Singhal et al., 2023a; Liu et al., 2023; Delsoz et al., 2023; Yang et al., 2022). Some studies discuss the application of LLMs in the field of medical education (Sallam, 2023; Safranek et al., 2023). Others have explored the challenges of practical clinical application of LLMs (Harrer, 2023; Singhal et al., 2023a; Sezgin, 2023). Med-PaLM2 is Google’s medical domain LLMs is the first LLM that can achieve an “expert” level of performance on the MedQA dataset of US Medical Licensing Examination (USMLE2)-style questions, with an accuracy of over 85% (Singhal et al., 2023b).
However, in the field of public healthcare, there are currently few relevant studies. Some studies provide suggestions from the perspective of experts on how the public can use LLMs to obtain healthcare information (Biswas, 2023). However, they do not provide an in-depth understanding of the general public’s perspectives, so the proposals given are often from a technical perspective and lack a sociological perspective. In 2023, a study examined the benefits and challenges of Carecall, an LLM-based conversational robot developed in Korea, for public health applications (Jo et al., 2023). Carecall is aimed at socially isolated individuals, while popular LLMs such as ChatGPT can potentially help the mass public access healthcare information. So in this paper, we aim to explore how the public utilizes general-purpose LLMs for healthcare purposes. A recent study explored the public’s intention to use ChatGPT for self-diagnosis (Shahsavar et al., 2023), but the survey was relatively simple – it only investigated whether the public was willing to use ChatGPT for self-diagnosis, and ignored many other related issues. Instead, we aspired to provide a nuanced picture of users’ motivations, experiences, and challenges when using LLMs for healthcare.
@@ 2.2. How the Public Access Healthcare Information
The Internet has greatly changed the form of information transmission, and information can be transmitted at an unprecedented speed. Before 2010, the public mainly obtained healthcare information from the Internet through search engines (Berland et al., 2001; Spink et al., 2004; Eysenbach and Kohler, 2003; Gray et al., 2005). After 2010, the methods of obtaining health information gradually diversified. Although search engines were still the most popular channel (De Choudhury et al., 2014), the public also began to try different methods to obtain health information. Some people used social media such as Twitter (De Choudhury et al., 2014; Scanfeld et al., 2010), WeChat (Zhang et al., 2017; Hong et al., 2021) or Instagram (Chung et al., 2017) to obtain or share health information. Due to the shortage of medical resources, some online consultation platforms have emerged – patients communicate with doctors online through medical platforms  (Ding et al., 2020; Hao, 2015; Hao et al., 2017). With the rise of artificial intelligence technology, some AI-driven medical/healthcare chatbots have emerged in recent years  (Pereira and Díaz, 2019), such as Ada  (You and Gui, 2020). Patients or the general public can enter questions into the chatbots and get answers. Overall, the ways for the public to obtain health information are becoming increasingly diverse, which is in line with the trend of “information explosion” in the information society. Diverse methods of obtaining health information can be likened to ”ecological niches” (Zhu et al., 2014) in the digital ecosystem (Milton et al., 2024; McLuhan, 1994), that is, like online communities, there is a competitive and complementary relationship between different methods of obtaining information. Previous studies have shown that there is heterogeneity in the acquisition of health information by search engines and social media, that is, users may search for different types of health information through search engines and social media (De Choudhury et al., 2014; Kanthawala et al., 2016), and the two methods are complementary.
LLMs demonstrate more powerful natural language interaction capabilities and have penetrated the medical field (Li et al., 2023). After LLMs join the ecological niche, which methods will compete with them and which will complement them remain under investigation.
2.3. Concerns When Accessing Healthcare Information using LLMs
@@
In the information explosion process, one of the prevalent concerns is that misinformation also explodes (Zhou et al., 2019; Shu et al., 2017). In the field of public health, misinformation is particularly rampant (Suarez-Lledo and Alvarez-Galvez, 2021). Misinformation related to serious diseases (e.g., cancer-related misinformation (Johnson et al., 2022; Chen et al., 2018; Warner et al., 2022)) has often been circulated on social media, accompanied by advertisements for high-priced health products that had no actual therapeutic effect. Since the beginning of the COVID-19 pandemic, conspiracy theories have emerged – for example, rumor went that the COVID-19 virus was spread by 5G technology (Ahmed et al., 2020). A wide range of misinformation about the COVID-19 pandemic existed on social media (Enders et al., 2020; Kouzy et al., 2020; Evanega et al., 2020; Cuan-Baltazar et al., 2020). Even more so, during the COVID-19 pandemic, it is not just the general public who may be spreading misinformation, but even professional doctors. One study identified 52 American doctors spreading misinformation on at least 5 different social platforms during the pandemic, which may have impacted millions of people (Sule et al., 2023).
@@
Existing studies have pointed out that ChatGPT could provide accurate health information if appropriately prompted (Mehnen et al., 2023). This proves the huge potential of large language models in disease diagnosis. However, considering that LLMs are known to generate “hallucinations” (Alkaissi and McFarlane, 2023; Rawte et al., 2023), understanding how the public experiences “hallucinations” when using LLMs to access healthcare information is important. One study showed that in a Turing test, patients had difficulty distinguishing whether the advice was given by a real doctor or generated by ChatGPT (Nov et al., 2023). Many doctors and scholars have warned that “AI cannot replace doctors (Mehnen et al., 2023; Li, 2023; Cheng et al., 2023).” Some researchers thought that although ChatGPT could provide useful information and answer general questions related to diseases, it cannot replace the expertise, training, and experience of a qualified doctor (Cheng et al., 2023). There is even a report of life-threatening delays in diagnosis due to the use of Chatgpt (Saenger et al., 2024).
@@
The doctor-patient relationship has always been a controversial topic, which can be complicated by LLMs. As early as 2006, there was research on doctors using search engines to diagnose conditions (Tang and Ng, 2006). In 2011, a survey showed that nearly half of doctors used Google and Yahoo to diagnose patients (Unknown, 2011). In November 2023, a video on the Chinese Internet sparked heated discussions. A doctor diagnosed a patient while swiping the screen of his mobile phone to search. Many Internet users believed that the doctor’s behavior was inappropriate (Unknown, 2023). There has also been a lot of discussion on social media platforms about whether doctors should use search engines to diagnose patients (None, 2017a). A relevant question is “whether doctors get mad when some patients come with their own Google research and diagnosis” (None, 2017b). These are typical issues and dilemmas in the doctor-patient relationship. Nowadays, the rise of LLMs and their domain knowledge in healthcare and medicine may have an extra impact on the doctor-patient relationship in addition to search engines and online communities. Will patients mind if doctors use LLMs to diagnose them? This question is worth further consideration and research.
The above contents can be summarized as the ethics of ChatGPT and other large language models in public healthcare. A recent study (Haltaufderheide and Ranisch, 2024) found that the existing literature, which is about the ethics of ChatGPT in medicine and healthcare, mainly focuses on clinical applications and the support of health professionals and researchers. But the studies on public health are rare.
@@ 3. Methodology
In this IRB-approved study, we used a combination of interviews and surveys to approach our research questions. Anyone over 18 who had used LLMs for healthcare Q&A or information search was eligible for our study. The two methods are complementary to each other: surveys were administered to understand users’ overall perceptions and ensure the generalizability of the findings. Specifically, through sampling statistics, we hope to understand the public’s participation and attitudes. For example, what proportion of the public uses LLMs to obtain health information; what proportion of the public supports doctors using LLMs for diagnosis, etc. Interviews were used to learn about users’ nuanced motivations, experiences, and concerns when using LLMs for health information searches.
@@ 3.1. Surveys
Between December 2023 and March 2024, we distributed the online survey on social media platforms such as X, Reedit, and Discord. In the recruitment message, we invited users who satisfied the inclusion criteria to complete our survey. The participation was voluntary, and the participants were offered a chance to enter a prize draw worth 50 dollars (for 5 participants). Since our advertising was primarily distributed through online communities, people who participated in the survey may be limited to those who were familiar with social media.
@@
Interested participants were directed to the survey after clicking on our Qualtrics link. They were first shown information related to the survey; after they gave consent, they proceeded to the survey questions. The survey had a total of 23 questions (Including three short answer questions) and the estimated completion time was 10 minutes – in the end, most participants completed the survey within an average of 5 minutes. When designing survey questions, we referred to previous research on AI healthcare chatbots (Nadarzynski et al., 2019) and took interview results and actual conditions into account. Specifically, the survey was divided into three parts: (1) Demographic information such as gender, ethnicity, education levels, and so on; (2) How participants searched for healthcare information and what types of healthcare information they searched for – for example, “Have you ever used the search engine (like Google, Bing, Baidu) to search for health information?”, “Have you ever searched for information about medication (treatment, drugs, hospitalization)?” (3) Participants were asked to evaluate controversial issues in the use of LLMs using a 5-point Likert scale(From “Strongly Disagree” to “Strongly Agree”), e.g, “I think it’s reasonable for doctors to use LLMs to diagnose and treat conditions.”
@@
During the quantitative analysis, we first removed surveys that were duplicates or looked like robot responses. When a large number of consecutive surveys appeared within a short period, we determined them as filled in by robots. We ended up receiving 219 valid surveys. Only 5 of these people had never used the Internet to search for healthcare information, so we finally used the remaining 214 responses for analysis. Similar to previous studies, we transformed most of the questions into binary variables and excluded neutral values (Nadarzynski et al., 2019). For ethnicity and age, we conducted population statistics by interval. The questions involving value judgments and personal opinions were complex and did not fully conform to the binary distribution. However, they could be used to explore the acceptability of LLMs in public healthcare applications and provide important guidance for future directions.
Among the 214 participants, most participants were aged 20-29 (100) and 30-39 (86). There were 20 participants aged 40-49 years. Only 6 participants were 10-19 years old and 3 participants were 50-59 years old. There were no participants in other age groups. This shows that most of the survey participants are young people. This age distribution is similar to the age distribution of Chatgpt users(None, 2017c). There were 111 males and 99 females, and 4 participants did not want to say their gender or were of a third gender. There is no statistical difference in the gender ratio. The largest number of participants were white (90), followed by Asians (57) and blacks or African Americans (51), with fewer other ethnic groups. The educational level of most participants is postsecondary (188).